# Preference Poisoning Attacks on Reward Model Learning

### v1

@@ 5.3. Experiment Setup
Our experiments involve four problem settings. In the first two, reward model learning is done in the context of value alignment for control (Christiano et al., 2017). The first of these generates control trajectories in MuJoCo environment associated with low-dimensional state inputs, whereas the second uses Atari to generate control trajectories with images as inputs. The third considers a recommendation system context using textual information. For this we use the Amazon rating data, [^1] https://cseweb.ucsd.edu/~jmcauley/datasets/amazon_v2/, Home and Kitchen. which we transform into pairwise comparisons. Our final experiment involves training a reward model toward safety alignment of Large Language Models (LLMs). We conducted 5 independent runs for each experiment, which means the data were collected independently for each run, and we report the mean and standard error of the mean (SEM) in the plots. The only exception is for LLM, where we conducted only 1 run due to the availability of the dataset. In most cases below, we learn a neural network reward model $R_{\theta}$, with one exception where we also investigate robustness of a linear reward model in comparison with a neural network in the same setting. In all cases, we vary the budget $B$ between 1% and 10% of data. Next we provide further details about the experiment environments and our setup; further details are provided in the Appendix.
MuJoCo Control: we use three MuJoCo environments: Reacher, Hopper, and Walker2D. We use trajectory data as inputs (outcomes), where trajectory $\mathcal{T}=((s_{1},a_{1}),(s_{2},a_{2}),\cdots(s_{l},a_{l}))$ is induced by a random policy $a_{i}=\pi(s_{i})$, with $s_{i}$ representing the system state (low-dimensional and observable) and $a_{i}$ the action. Following Christiano et al. (2017), we record the total reward returned by the MuJoCo simulator as the reward for the trajectory. The trajectory with the higher reward is marked as winning, and the other as losing. The trajectory length for each environment is 30.
Atari Vision-Based Control: For Atari control with image inputs, we experimented with three environments: Pong, Breakout, and Qbert. We used trajectory data as inputs as in the previous setting, except now our states $s_{i}$ are images. We concatenate actions with image embeddings in the mid-linear layer. For experiments with PCA we reduces the input dimension to $20$, and we then concatenate it with the action.
Recommendation System: We used Amazon Rating Data collected with the purpose of designing recommendation systems. While this, like many conventional recommendation datasets, relies on user ratings, it has been observed that learning utilities using pairwise comparisons rather then relying solely on ratings can yield better recommendations (Kalloori et al., 2018). Consequently, we used this data to generate a derived dataset of pairwise comparisons by comparing the recorded ratings of pairs of comments, which were used as inputs (outcomes). We used BERT to generate embeddings for each input comment, considering it as the input representation. When we apply PCA instead, we reduce the original dimension to 20 components. In this domain, we experiment using both a linear neural network and a 3-layer neural network with ReLU activations (multilayer perceptron (MLP)). For linear neural networks, the average training accuracy was $92.3\%$, and the average test accuracy was $91.5\%$. For the neural network architecutre, average training accuracy was 100%, while average test accuracy was 92.3%. While it seems that the latter architecture is superior to the linear model, our experiments below offer an interesting qualification in that regard.
@@
Safety Alignment: Finally, we used the PKU-SafeRLHF-30k dataset [^2] https://huggingface.co/datasets/PKU-Alignment/PKU-SafeRLHF-30K, which includes 30k question-answer pairs annotated with both helpfulness and harmlessness preference labels. For the scope of our study, we only use the harmlessness preference labels for our safety alignment task. Then we train the reward model initialized with an instruction following language model, which is obtained by performing supervised fine-tuning on the pre-trained model LLaMA-7B (Touvron et al., 2023) using the Stanford Alpaca Dataset.
5.4. Results
@@
MuJoCo Control
Figure 2. Promotion attack efficacy in MuJoCo. Top row: success rate. Bottom row: accuracy.
We begin by considering promotion attacks in the low-dimensional MuJoCo setting. Figure 2 (top) presents attack success rate for the best-performing variants of each attack class (gradient-based, RBD), as well as the baseline (random) which flips comparison labels of randomly selected datapoints. Our first observation, which will be echoed in other domains below, is that attack success rate (and, thus, vulnerability) varies quite significantly with problem setting. For example, success rate in Hopper reaches over 60% with 5% of the attack budget, and approximately 90% with 10% of the budget. On the other hand, success rate for Reacher and Walker is just over 50% even with 10% of the budget.
Our second observation is that in all three domains, the (best) gradient-based method is most effective, whereas MLE is quite robust to random label flips. Nevertheless, we do note that the (best) RBD heuristic is typically competitive with the best attack overall, particularly at small budgets. For example, at 1-5% of the budget, the performance of both classes of attacks is similar.
@@
Figure 3. Relative efficacy of gradient-based attacks (top row) and RBD attacks (bottom row) in MuJoCo.
In Figure 3, we consider now ablations in terms of different gradient-based approaches (top) and RBD methods (bottom). Here we make two observations. First, in this setting, pretraining the neural network before performing the gradient-based attack, as done by Koh and Liang (2017), performs extremely poorly, yielding a nearly zero success rate. This is likely because it is fragile to uncertainty about neural network initialization. Rather, our primary approach in which we train the attack to be explicitly robust to initialization uncertainty performs best. In addition, we note that black-box attacks in this context are essentially as effective as white-box attacks, suggesting that we do not need precise knowledge of the neural network architecture. Second, we generally find that RBD-Norm outperforms RBD-Reward in two of the three environments, with the two exhibiting similar performance in the third (Hopper).
Figure 4. Demotion attack efficacy in MuJoCo.
@@
Atari Vision-Based Control
Next, we turn to the more challenging control setting in which true state of the controlled system is not directly observable, but is instead observed using (relatively) high-dimensional visual perception. This, in turn makes gradient-based attacks significantly more difficult to scale.
Figure 5. Promotion attack efficacy in Atari. Top row: success rate. Bottom row: accuracy.
@@
Figure 6. Relative efficacy of gradient-based attacks (top row) and RBD attacks (bottom row) in Atari.
Next, we consider again the relative efficacy of gradient-based methods (Figure 6, top row). In this setting, surprisingly, the most effective attack is to first preprocess the high-dimensional input images using PCA (obtaining 20 dimensions), and perform a (black-box) attack on a neural network over this reduced-dimensional outcome space. This approach is significantly better than making use of the low-dimensional embedding learned as part of the reward model $R_{\theta}$ trained on clean data. Additionally, making use of the conjugate-gradient method to approximate the inverse in implicit gradient computation, as suggested by Koh and Liang (2017) was extremely slow, with a single attack taking $\sim$ 12 hours for one seed.
@@
Our observation that relative performance of different attack methods varies considerably by domain and environment is instructive: our consideration of two classes of attack algorithms, and multiple variants in each class thereby demonstrates the value of comprehensive vulnerability analysis that this provides. This is in contrast to common vulnerability analysis of ML to poisoning attacks in prior work, where only a single attack algorithm is typically evaluated (e.g., (Jha et al., 2023; Xiao et al., 2012; Zhang et al., 2020b)).
Figure 7. Demotion attack efficacy in Atari.
Finally, we consider demotion attacks in the Atari control domain. The results are shown in Figure 7, are are broadly similar to what we have observed for promotion attacks. However, here we see that RBD heuristics tend to outperform the best gradient-based method in nearly all cases. This may seem surprising at first, given that the more expensive gradient-based methods are generally viewed as highly principled, but note that these are also heuristic in a number of ways. For example, in this discrete setting, gradient-based approaches rely on relaxation of the discrete decision variables. Moreover, as the basic variant of these fails to scale to the higher-input-dimension problem posed in vision-based control, we now also rely on the linear (PCA) dimensionality reduction technique blended with the gradient-based method (which does outperform a non-linear learned embedding). In any case, this again reinforces our broader observation that the nature of the best attack can vary by domain, environment, and—in this case—attacker’s objective.
Recommendation System
Next, we shift away from control problems, where reward modeling learning is motivated by (and is a key step for) value alignment, to a recommendation system setting. As mentioned above, we postprocess an existing recommendation rating dataset to pairwise comparison data, which we then poison.
Figure 8. Efficacy of promotion (left) and demotion (right) attacks on the recommendation dataset (neural network model).
In Figure 8 we show the results of poisoning in this context for both promotion and demotion attacks, when we learn a neural network reward model (MLP). Notably, in this setting, which is also too high-dimensional for a direct application of the gradient-based approach, RBD methods are significantly more effective than gradient-based algorithms, both in the case of promotion and demotion attacks. Additionally, the best method (RBD) in this setting achieves 100% success rate in both types of attacks with only a 5% attack budget, and 60-70% success rate with a mere 1% attack budget.
Figure 9. Accuracy of promotion and demotion attacks on the recommendation dataset (neural network model).
@@
Figure 10. Relative efficacy of gradient-based attacks (top) and RBD attacks (bottom) on the recommendation dataset (neural network model).
@@
Figure 11. Efficacy of promotion and demotion attacks on the recommendation dataset (linear model).
Thus far, our consideration of reward model poisoning was focused on neural network reward models. We now explore the extent to which using linear models instead—which are inferior in conventional terms, such as prediction accuracy—yields better or worse robustness to poisoning attacks.
Figure 12. Accuracy of promotion and demotion attacks on the recommendation dataset (linear model).
@@
Figure 13. Relative efficacy of gradient-based attacks (top) and RBD attacks (bottom) on the recommendation dataset (linear model).
Finally, Figure 13 presents the ablation results comparing different gradient (top) and RBD (bottom) approaches in terms of relative efficacy. In this domain, we find an interesting difference from our observations so far: using a pre-trained model as part of the gradient-based scheme now outperforms the approach that instead using $K$ random initializations, especially in the demotion attack. In RBD comparison, on the other hand, RBD-Norm handily outperforms RBD-Reward in this setting, comparable to our observations above.
Safety Alignment
Our final consideration involves the safety alignment problem in the context of language models. Specifically, we consider a dataset of pairwise comparisons involving relative harmfulness of prompt responses. We learn both the clean, and poisoned reward model $R_{\theta}$ by fine-tuning the LLaMA-7B model using the Stanford Alpaca Dataset.As gradient-based methods are not scalable in this setting due to the size of the neural networks, we only consider the RBD-based approaches for attacks.
Figure 14. Efficacy of promotion (left) and demotion (right) attacks in terms of success rate (top) and stealth, as measured by test accuracy (bottom) in the LLM safety alignment setting.
Figure 14 presents the results on the efficacy of the poisoning attacks (top) and accuracy degradation (bottom). What we observe is that the LLM setting is particularly vulnerable, far more than all of the previous settings we consider. Specifically, the best attack achieves 100% success rate for both promotion and demotion goals with only 0.3% poisoned instances. A key reason is the structure of this and similar datasets: outcome comparisons are made with respect to a relatively limited set of prompts and responses, which entails many identical outcomes in pairwise comparisons. Consequently, it is often the case that there are a large set of outcomes similar to the target, and flipping the comparison label for all of these (or any subset) is highly efficacious. We also observe that in this setting, we see that both RBD-Reward and RBD-Embedding (using Llama embedding) perform comparably. In addition, we again see minimal accuracy degradation as a result of the attack.
6. Effectiveness of Defense

### v2

@@ 4.2 Rank-by-Distance Approaches
4.3 Theoretical Analysis
Recall that an important theoretical question in our proposed approach is whether focusing on the finite-sample approximation in Problem (3) is principled, in the sense that it yields a provable approximation guarantee even vis-a-vis the true objective (Problem (2)). We now address this issue, focusing on promotion attacks; the analysis for demotion attacks is essentially identical. Note that the answer is not self-evident. First, we are not merely approximating the expectation with sample average (and removing the constant factor, which does not impact the optimal solution), but approximating an optimization problem. Thus, law of large numbers does not immediately imply that the quality of solution to the approximate problem converges. Second, the problem involves bi-level optimization, since the attacker’s optimal solution depends, in turn, on the learning optimization problem that yields the reward model parameters $\theta$. Next, we prove that this problem has polynomial sample complexity, that is, the number of samples $N$ grows only polynomially in $1/\epsilon$ and the dimension of the function class $R_{\theta}$, where $\epsilon$ represents the quality of the approximation of the solution to Problem (2) by the solution to Problem (3).
Lemma 4.1.
Suppose $\mathcal{F}$ is a vector space of real-valued functions, and $H=\{\mathbf{1}_{\{f+f(a)\geq 0\}}:f\in\mathcal{F}\}$, where $a$ is a constant, then VCdim$(H)\leq\text{dim}(\mathcal{F})+1$, where $\text{dim}(\mathcal{F})$ is the dimension of the vector space of functions $\mathcal{F}$.
Proof.
Theorem 4.2 ([1]).
Here $m_{L}(\varepsilon,\gamma)=\min\{m:m$ is a sufficient sample size for $(\varepsilon,\gamma)$-learning $H\text{ by }L\},$ meaning for $m\geq m_{0}(\varepsilon,\gamma)$ and $z\in Z^{m}$ chosen according to $P^{m}$, $P^{m}\{\text{er}_{P}(L(z))\geq\text{opt}_{P}(H)-\varepsilon\}\geq 1-\gamma.$
Theorem 4.3.
Let =$\mathcal{F}=\{R_{\theta}|\theta\in\Theta\}$ and suppose $\text{dim}(\mathcal{F})=d$. Then for all $\epsilon>0$, and for all $N>m_{0}(\epsilon,\gamma)$ where $m_{0}(\epsilon,\gamma)=\frac{64}{\varepsilon^{2}}\left(2(d+1)\ln\left(\frac{12}{\varepsilon}\right)+\ln\left(\frac{4}{\gamma}\right)\right)$, $\max_{\delta\in\Omega}\hat{F}(\delta)\geq\max_{\delta\in\Omega}{F}(\delta)-\varepsilon$ with probability at least $1-\gamma$.
Proof.
Similarly, Problem (3) is equivalent (up to $1/N$ factor) to
5 Experiments
@@ 5.1 Experiment Setup
Our experiments involve four problem settings. Our first experiment involves training a reward model toward safety alignment of Large Language Models (LLMs). In the next two, reward model learning is done in the context of value alignment for control [16]. The first of these generates control trajectories in MuJoCo environment associated with low-dimensional state inputs, whereas the second uses Atari to generate control trajectories with images as inputs. The fourth considers a recommendation system context using textual information. For this we use the Amazon rating data,111cseweb.ucsd.edu/~jmcauley/datasets/amazon_v2/, Home and Kitchen. which we transform into pairwise comparisons. The motivation for exploring these four settings is to assess the performance and impact of our proposed attack for qualitatively different control problems, since the preference-based reward model learning and RLHF have been used in domains as distinct as robotic control [35] and LLMs [62].
We conducted 5 independent runs for all but one experiment, which means that the data were collected independently for each run, and report the mean and standard error of the mean (SEM) in the plots. The only exception is LLM, where we conducted only 1 run due to high computational requirements. In most cases below, we learn a neural network reward model $R_{\theta}$, with one exception where we also investigate robustness of a linear reward model in comparison with a neural network in the same setting. In all cases, we vary the budget $B$ between 1% and 10% of data. Next we provide further details about the experiment environments and our setup; further details are provided in the Appendix.
@@
Safety Alignment: We use the PKU-SafeRLHF-30k dataset222huggingface.co/datasets/PKU-Alignment/PKU-SafeRLHF-30K, which includes 30k question-answer pairs annotated with both helpfulness and harmlessness preference labels. For the scope of our study, we only use the harmlessness preference labels for our safety alignment task. Then we train the reward model initialized with an instruction following language model, which is obtained by performing supervised fine-tuning on the pre-trained model LLaMA-7B [83] using the Stanford Alpaca Dataset.
MuJoCo Control: We use three MuJoCo environments: Reacher, Hopper, and Walker2D. We use trajectory data as inputs (outcomes), where trajectory $\mathcal{T}=((s_{1},a_{1}),(s_{2},a_{2}),\cdots(s_{l},a_{l}))$ is induced by a random policy $a_{i}=\pi(s_{i})$, with $s_{i}$ representing the system state (low-dimensional and observable) and $a_{i}$ the action. Following [16], we record the total reward returned by the MuJoCo simulator as the reward for the trajectory. The trajectory with the higher reward is marked as winning, and the other as losing. The trajectory length for each environment is 30.
Atari Vision-Based Control: For Atari control with image inputs, we experiment with three environments: Pong, Breakout, and Qbert. We used trajectory data as inputs as in the previous setting, except now our states $s_{i}$ are images. We concatenate actions with image embeddings in the mid-linear layer. For experiments with PCA we reduces the input dimension to $20$, and we then concatenate it with the action.
Recommendation System: Finally, we use Amazon Rating Data collected with the purpose of designing recommendation systems. While this, like many conventional recommendation datasets, relies on user ratings, it has been observed that learning utilities using pairwise comparisons rather then relying solely on ratings can yield better recommendations [43]. Consequently, we used this data to generate a derived dataset of pairwise comparisons by comparing the recorded ratings of pairs of comments, which were used as inputs (outcomes). We used BERT to generate embeddings for each input comment, considering it as the input representation. When we apply PCA instead, we reduce the original dimension to 20 components. In this domain, we experiment using both a linear neural network and a 3-layer neural network with ReLU activations (multilayer perceptron (MLP)). For linear neural networks, the average training accuracy was $92.3\%$, and the average test accuracy was $91.5\%$. For the neural network architecutre, average training accuracy was 100%, while average test accuracy was 92.3%. While it seems that the latter architecture is superior to the linear model, our experiments below offer an interesting qualification in that regard.
5.2 Results
@@
Safety Alignment. Our first consideration involves the safety alignment problem in the context of language models. Specifically, we consider a dataset of pairwise comparisons involving relative harmfulness of prompt responses. We learn both the clean, and poisoned reward model $R_{\theta}$ by fine-tuning the LLaMA-7B model using the Stanford Alpaca Dataset. As gradient-based methods are not scalable in this setting due to the size of the neural networks, we only consider the RBD-based approaches for attacks.
Figure 2: Efficacy of promotion (left) and demotion (right) attacks in terms of success rate (top) and stealth, as measured by test accuracy (bottom) in the LLM safety alignment setting.
Figure 2 presents the results on the efficacy of the poisoning attacks (top) and accuracy degradation (bottom). What we observe is that the LLM setting is extremely vulnerable: the best attack achieves 100% success rate for both promotion and demotion goals with only 0.3% poisoned instances. A key reason is the structure of this and similar datasets: outcome comparisons are made with respect to a relatively limited set of prompts and responses, which entails many identical outcomes in pairwise comparisons. Consequently, it is often the case that there are a large set of outcomes similar to the target, and flipping the comparison label for all of these (or any subset) is highly efficacious. We also observe that in this setting, both RBD-Reward and RBD-Embedding (using Llama embedding) perform comparably. In addition, we see minimal accuracy degradation as a result of the attack.
Figure 3: Promotion attack efficacy in MuJoCo. Top row: success rate. Bottom row: accuracy.
Next, we evaluate the downstream impact of reward model attacks on RLHF in this setting. The success rates (of the learned policy choosing the target response to a prompt) are provided in Table I for a randomly chosen target candidate, for whom baseline success rate (with no attack) is $<1$%.
TABLE I: Promotion attack efficacy in RLHF.
Poisoning Ratio (%) 1 3 5
Attack Success Rate (%) 93 93.5 99
What we can see is that while the efficacy of the attack degrades somewhat compared to targeting reward model learning directly, it is nevertheless quite successful, achieving 99% success rate with only 5% of labels flipped.
MuJoCo Control. Next we consider promotion attacks in the low-dimensional MuJoCo setting. Figure 3 (top) presents attack success rate for the best-performing variants of each attack class. Our first observation, which will be echoed in other domains below, is that attack success rate (and, thus, vulnerability) varies quite significantly with problem setting. For example, success rate in Hopper reaches over 60% with 5% of the attack budget, and approximately 90% with 10% of the budget. On the other hand, success rate for Reacher and Walker is just over 50% even with 10% of the budget.
Our second observation is that in all three domains, the (best) gradient-based method is most effective. Nevertheless, we do note that the (best) RBD heuristic is typically competitive with the best attack overall, particularly at small budgets. For example, at 1-5% of the budget, the performance of both classes of attacks is similar.
@@
Figure 4: Relative efficacy of gradient-based attacks (top row) and RBD attacks (bottom row) in MuJoCo.
In Figure 4, we consider now ablations in terms of different gradient-based approaches (top) and RBD methods (bottom). Here we make two observations. First, in this setting, pretraining the neural network before performing the gradient-based attack, as done by [44], performs extremely poorly, yielding a nearly zero success rate. This is likely because it is fragile to uncertainty about neural network initialization. Rather, our primary approach in which we train the attack to be explicitly robust to initialization uncertainty performs best. In addition, we note that black-box attacks in this context are essentially as effective as white-box attacks, suggesting that we do not need precise knowledge of the neural network architecture. Second, we generally find that RBD-Norm outperforms RBD-Reward in two of the three environments, with the two exhibiting similar performance in the third (Hopper).
Figure 5: Demotion attack efficacy in MuJoCo.
@@
Atari Vision-Based Control. Next, we turn to the more challenging control setting in which true state of the controlled system is not directly observable, but is instead observed using (relatively) high-dimensional visual perception.
Figure 6: Promotion attack efficacy in Atari. Top row: success rate. Bottom row: accuracy.
@@
Figure 7: Relative efficacy of gradient-based attacks (top row) and RBD attacks (bottom row) in Atari.
Next, we consider again the relative efficacy of gradient-based methods (Figure 7, top row). In this setting, surprisingly, the most effective attack is to first preprocess the high-dimensional input images using PCA (20 dimensions), and perform a (black-box) attack on a neural network over this reduced-dimensional outcome space. This approach is significantly better than making use of the low-dimensional embedding learned as part of the reward model $R_{\theta}$ trained on clean data. Additionally, making use of the conjugate-gradient method to approximate the inverse in implicit gradient computation, as suggested by [44] was extremely slow, with a single attack taking $\sim$ 12 hours for one seed.
@@
Our observation that relative performance of different attack methods varies considerably by domain and environment is instructive: our consideration of two classes of attack algorithms, and multiple variants in each class thereby demonstrates the value of comprehensive vulnerability analysis that this provides. This is in contrast to common vulnerability analysis of ML to poisoning attacks in prior work, where only a single attack algorithm is typically evaluated (e.g., [39, 99, 108]).
Figure 8: Demotion attack efficacy in Atari.
Finally, we consider demotion attacks in the Atari control domain. The results are shown in Figure 8, are are broadly similar to what we have observed for promotion attacks. However, here we see that RBD heuristics tend to outperform the best gradient-based method in nearly all cases. This may seem surprising at first, given that the more expensive gradient-based methods are generally viewed as highly principled, but note that these are also heuristic in a number of ways. For example, in this discrete setting, gradient-based approaches rely on relaxation of the discrete decision variables. Moreover, as the basic variant of these fails to scale to the higher-input-dimension problem posed in vision-based control, we now also rely on the linear (PCA) dimensionality reduction technique blended with the gradient-based method (which does outperform a non-linear learned embedding). In any case, this again reinforces our broader observation that the nature of the best attack can vary by domain, environment, and attacker’s objective.
Figure 9: Efficacy of promotion (left) and demotion (right) attacks on the recommendation dataset (neural network model).
Recommendation System. Finally, we consider a recommendation system setting. In Figure 9 we show the results of poisoning in this context for both promotion and demotion attacks, when we learn a neural network reward model (MLP). Notably, in this setting, which is also too high-dimensional for a direct application of the gradient-based approach, RBD methods are significantly more effective than gradient-based algorithms, both in the case of promotion and demotion attacks. Additionally, the best method (RBD) in this setting achieves 100% success rate in both types of attacks with only a 5% attack budget, and 60-70% success rate with a mere 1% attack budget.
Figure 10: Accuracy of promotion and demotion attacks on the recommendation dataset (neural network model).
@@
Figure 11: Relative efficacy of gradient-based attacks (top) and RBD attacks (bottom) on the recommendation dataset (neural network model).
@@
Figure 12: Efficacy of promotion and demotion attacks on the recommendation dataset (linear model).
Thus far, our consideration of reward model poisoning was focused on neural network reward models. We now explore the extent to which using linear models yields better or worse robustness to poisoning attacks.
Figure 13: Accuracy of promotion and demotion attacks on the recommendation dataset (linear model).
@@
Figure 14: Relative efficacy of gradient-based attacks (top) and RBD attacks (bottom) on the recommendation dataset (linear model).
Finally, Figure 14 presents ablation results comparing different gradient (top) and RBD (bottom) approaches in terms of relative efficacy. In this domain, we find that using a pre-trained model as part of the gradient-based scheme now outperforms the approach that instead uses $K$ random initializations, especially in the demotion attack. In RBD comparison, on the other hand, RBD-Norm handily outperforms RBD-Reward in this setting, comparable to our observations above.
Partial Knowledge of Data. Our experiments thus far have assumed that the attacker observes the entire dataset $D$. We now consider the impact of relaxing this assumption, when only partial information about $D$ is available. The results are provided in Figure 15 for the MuJoCo and Atari control environments, and demonstrate that while attack success rate degrades under partial knowledge of data, it does so relatively slowly.
Figure 15: Promotion attack with partial knowledge of data (best attack results). MuJoCo (top) and Atari (bottom).
6 Effectiveness of Defense